# On the Structure and Semantics of Identifier Names Containing Closed Syntactic Category Words

### v1

@@ 4.2 Manual Process for Annotating Part-of-Speech
Initially, one author (annotator) is assigned annotate each identifier in the CCID with its grammar pattern. The annotator has experience annotating identifiers with PoS from prior work Newman_GP ; peruma_test_name_gp . The process is as follows: The annotator is given a split (using Spiral HuckaSpiral ) identifier along with the identifier’s type, file path, and line number to make it easy to find the identifier in the original code. The annotator is allowed to look at the source code from which the identifier originated if necessary. The annotator is asked to additionally identify and correct mistakes made by Spiral. When the annotator is finished, two additional annotators are asked to validate (agree or disagree) with the annotations created by the original annotator. Any disagreements are discussed and fixed, if required. Furthermore, a fourth annotator made annotations, which are then compared to the original annotator’s work. Again, disagreements are discussed and fixed. An example disagreement is with the identifier where_len, which is a tricky one because ‘where’ is typically an adverb or conjunction. However, in this case ‘where’ is a reference to a void pointer variable called ‘where’ within the code. So ‘where_len’ is the length of the memory this pointer points to, making ‘where’ a noun-adjunct in this case; describing the type of length. Thus, its grammar pattern is NM N. The Fleiss’ Kappa for this process was .916.
We did not expand abbreviations for a couple reasons. The first is that some abbreviations are more meaningful than their expanded terms (e.g., HTTP, IPV4, SSL) due to how frequently they are used in their abbreviated form by the community. The second reason is that abbreviation expansion techniques are not widely available and vary widely in terms of effectiveness on different types of terms newmanabbrev ; Jingxuan:2023 . Therefore, a realistic worst-case scenario for developers and researchers is that no abbreviation-expansion technique is available to use, and their PoS taggers must work in this worst-case scenario. Whenever we recognized one, we do not split domain-term abbreviations (e.g., Spiral will make IPV4 into IPV 4; we corrected this to IPV4). We do this because it is the view of the authors that they should be recognized and appropriately tagged in their abbreviated (i.e., their most common) form.
@@
Determiner Conjunction Preposition Digit
@@ 5 Evaluation of RQ1: What behavioral roles do closed-category terms play in source code identifiers?
Our evaluation aims to establish, through RQ1 and RQ2: 1) how closed-category terms are used to convey differing types of program behavior, 2) the typical grammatical structure of identifiers containing closed-category terms, and 3) how closed-category term distributions differ across programming context, language, and system domains. This research question investigates the semantic role of closed-category grammatical patterns in identifier naming. We focus on four closed-category part-of-speech types: prepositions, digits, determiners, and conjunctions. We present our findings by (1) describing each category’s semantic function using axial codes, (2) summarizing behavioral trends via selective coding, and (3) highlighting shared trends through cross-category synthesis.
5.1 Methodology: Manual Process for Grounded Theory Annotations
We employed a grounded theory approach to analyze variable names in source code and their relationship to program behavior. This multi-phase coding process involved four annotators and combined individual annotation with iterative validation and synthesis to construct a theory grounded in observed naming patterns. The sample used in this study is a subset of the CCID described in Section 4. To construct this subset, we took the top 10 most common grammar patterns (Table 7) and collected all identifiers that followed these patterns; randomly selecting the 10th grammar pattern if multiple patterns had the same frequency. These represent the most common names (i.e., from the perspective of grammatical structure) used in the data set. This totals to 618 identifiers.
@@ 5.1 Methodology: Manual Process for Grounded Theory Annotations
Annotators grouped similar open codes into higher-level axial codes, focusing on patterns where particular grammar structures aligned consistently with specific behavioral roles. These axial codes captured mid-level abstractions (e.g., State Variables, Event Triggers), and were documented in the spreadsheet alongside notes justifying the grouping where needed. The Fleiss’ Kappa for this phase was: .971 for Digits, .996 for Determiners, .976 for Prepositions, and 1.0 for Conjunctions.
Each annotator’s axial codes were reviewed by a different annotator for validation. This cross-review process involved reading both the open codes and the proposed axial codes, discussing disagreements, and refining the categories until consensus was reached.
@@ 5.2 Digits in Identifiers
Digits in identifiers act as compact, semantic indicators of structure, ordering, or version. They are also often used to disambiguate entities and encode numeric conventions. Their meaning is typically inferred through domain knowledge, making them powerful, but potentially hard to understand for those without the requisite domain knowledge.
@@
We created a dual-axis framework for interpreting the meaning of digits, inspired by a single-axis framework we created in prior work on digits in identifiers peruma_digits . This framework reflects our observation that digits contribute information in two distinct ways: (1) the role they play within the local context (e.g., indexing, versioning), and (2) the source of meaning they draw from, which is often external to the immediate source code scope (e.g., domain conventions, technical standards). Every digit in the set has both a role and a source of meaning; they must be combined to fully understand the digit. We put an ‘x’ between each combination of ‘Role’ and ‘Source of Meaning’ Axial Code.
  • Role: What functional purpose the digit serves in the identifier.
    • Distinguisher: The digit differentiates conceptually similar entities, typically to avoid name collision errors from the compiler (e.g., arg1, tile2).
    • Version Identifier: The digit encodes versioning information such as protocol revisions or data format versions (e.g., http2, v1).
  • Source of Meaning: Where the interpretation of the digit originates, typically via convention, tooling, or domain-specific logic.
    • Auto-Generated: The digit is added automatically by tools, compilers, or naming systems to avoid conflicts (e.g., var1_2, jButton3).
    • Human-Named Convention: The digit’s meaning is primarily derived from ad hoc developer intent and is not more complex than distinguishing entities manually (e.g., str1, feature2).
    • Locally Specific Concept: The digit conveys project- or context-specific information, often related to coordinate systems, data structures, or memory layouts (e.g., m33 for matrix row 3 col 3).
    • Technology Term / Standard: The digit is part of a recognized domain-specific label, format, or protocol (e.g., HTTP2, Neo4j).
@@
    Description: This group captures identifiers where digits encode positional or logical roles based on system-specific conventions, such as grid layout or data structure indexing.
@@
    Description: This group captures identifiers that include digits as part of standardized or domain-specific naming conventions, often encoding formats or specifications.
@@
    Description: This group captures identifiers where the digit signals the version number of a protocol, tool, or technology component.
@@
    Description: This group captures identifiers that are automatically suffixed with a digit to ensure uniqueness, often generated by tools or compilers.
@@
Consider the identifier m34, which appears in the context of a matrix operation. To fully interpret the digit 3 in this name, we must consider both its Role and its Source of Meaning. Semantically, the digit serves as a Distinguisher; uniquely identifying this variable apart from its siblings (such as m32 and m31). However, its full interpretation depends on its Locally Specific Concept source: the developers have an internal convention that 3 refers to the row index, while 4 refers to the column. Without knowing the Source of Meaning, the numbers can only be interpreted as distinguishing one identifier from another; the meaning of the digits would remain ambiguous. This illustrates how both axes work together—Role tells us what the digit is doing, while Source of Meaning tells us how to interpret the value.
@@
Digits serve as semantic compression tools in source code: conveying versioning, layout, ordering, or configuration state using a minimal footprint. Their power lies in the shared assumptions between the name’s author and its reader. Whether distinguishing hosts (host1, host2), signaling protocol versions (http2), or denoting matrix dimensions (m33), digits rely on prior knowledge to be effective. This makes them:
@@
  • Hard to understand when overused without documentation or when the reader lacks background information/experience
Digits are structural shortcuts in the mental models of developers; a quick way to convey a lot of information in a small number of characters.
@@ 5.3 Prepositions in Identifiers
Prepositions in identifiers express spatial, temporal, or logical relationships. They are the most versatile (i.e., most axial codes) and frequently used closed-class grammatical structure in our dataset. Prepositions typically convey transformation, control conditions, event triggers, source origin, or context membership. Because only a subset of these are dual axis (Boolean Flow), we inline the definitions with our examples, unlike with Digits, where we separated them.
@@ 5.4 Determiners in Identifiers
    Definition: This group captures identifiers that reference a population or subset, typically using quantifiers like all, any, or some to guide iteration, filtering, or policy logic.
@@ 5.5 Conjunctions in Identifiers
Conjunction-based identifiers are rare but expressive. They signal compound behavior, dual-mode interfaces, or gated logic—often making hidden control flow or semantic relationships visible. Their rarity likely stems from the fact that developers often express conjunctions in logic rather than names. But when used, they highlight either an intent to foreground behavior or to capture structural duality within a single name.
@@ 5.6 Cross-Category Synthesis
Across digits, determiners, prepositions, and conjunctions, developers use closed-class grammatical structures to encode compact, behavior-rich semantics in identifiers. While each part-of-speech (POS) category exhibits distinct tendencies, analysis of grammar patterns reveals broader functional themes and stylistic consistencies across categories.
@@
It is interesting that booleans appear in all three of these contexts, but each is a different flavor; a way of expressing behavior that is unique to the closed-category terms used in the boolean identifier.
@@
    Digits imply procedural differentiation (method1, step2) or timeline indexing when appearing in coordinated identifiers (m31, m32).
@@
    Digits embed indexing, uniqueness, and domain roles
@@
These layered forms operate as semantic shortcuts to express complex behavior with very few words. They compress conditions, transformations, order, and relationships into concise forms.
@@ 5.7 Summary of RQ1
Through grounded theory analysis of closed-category terms in identifiers, we have uncovered and explored the ways in which these compact grammatical forms play a central role in expressing program behavior. Each part-of-speech category contributes distinct semantic functions, ranging from transformation and scoping to control flow and logical composition.
@@
Determiner Conjunction Digit Preposition
@@
We begin by analyzing the distribution of four closed categories: prepositions, determiners, conjunctions, and digits, across five programming contexts and three programming languages. We discuss which categories are most frequent in which contexts/languages and consider how those patterns may reflect the communicative goals of the developer. We then extend this analysis to system-level domain context, comparing the normalized frequency of closed-category term usage between domain-specific and general-purpose systems.
@@ 6.1 Closed-Category Term Usage Across Programming Contexts, Programming Languages, and System Domains
We now examine how differing contexts and closed-category grammar patterns relate to one another, and whether programming language further conditions their usage. We begin by analyzing cross-language correlations in closed-category term usage, followed by an exploration of correlations in how closed-category terms are used in different program contexts. We provide Table 8, which shows frequencies and percentages for PoS and terms, to help the reader understand what types of terms are most prevalent. However, for this RQ we rely primarily on Tables 9, 10, 11, and 12, which show the results of our Pearson Chi-squared tests and Standardized Pearson residuals. Using these, we highlight common patterns, terms, and the contexts or languages to which these patterns are correlated.
@@ 6.1.1 Language-Specific Differences in Closed-Category Term Usage
Methodology. To perform our Chi-Square test, we use the CCID described in Section 4.1. We count how many times each closed-category PoS appears in C++, Java, or C code by analyzing all 1,001 identifiers that contain closed-category terms. For example, we might find that there were 20 Digits in our data set found in C++ code, and 5 Digits in Java code. Once we have these frequencies, we apply the Chi-Square test, and Standardized Pearson residuals w/Bonferonni correction to determine overall significance, and per-part-of-speech significance, respectively.
@@ C C++ Java
  • Digits (D) are modestly underrepresented in Java (residual = –1.65), suggesting a mild tendency to avoid numeric suffixes in Java naming.
@@
While we did not find significant statistical evidence linking closed-category tag usage to programming language, the residual analysis and qualitative trends suggest mild idiomatic differences, particularly around digit usage and determiner phrasing. For example, in our Axial Code data from RQ1, Population/Subpopulation Reference identifiers were found in Java (21, 50%) and C++ (18, 42%) more than in C(3, 7%). These patterns may reflect broader stylistic conventions or design idioms of each language, but should be interpreted cautiously given the statistical outcome.
@@ 6.1.2 Context-Specific Differences in Closed-Category Term Usage
Methodology. To perform our Chi-Square test, we use the CCID described in Section 4.1. We analyzed all 1,001 identifiers that contained closed-category terms, and count how many times a closed-category term appears in one of our five code contexts: Attribute, Function, Class, Declaration, or Parameter. Once we have these frequencies, we apply the Chi-Square test, and Standardized Pearson residuals w/Bonferonni correction to determine overall significance, and per-part-of-speech significance, respectively.
@@
The selective coding data from RQ1 offers an explanation for this pattern. Conjunction-based grammar patterns tend to express compound logic, dual-purpose behavior, or guarded activation, which are most relevant when naming behaviors or actions rather than static values. For example, Guarded Action / Conditional Enablement patterns such as load_if_needed or activate_if_enabled appear in function names to encode preconditions or gating logic directly into the identifier.
@@
While we did find a significant correlation with Conjunctions and functions, it is important to recall that we have a limited number of Conjunctions in our dataset, at 16, meaning that while we have found potential trends, further research is required.
@@
Digits (D). Closed-category grammar patterns that include digits often feature numerals such as 1, 2, 0, 3, and 4, as shown in Table 8. Digits are significantly positively correlated with parameter names and class names, and significantly negatively correlated with function names (Standardized Pearson residual = 4.72 for class names, 4.25 for parameters, and –5.51 for functions; Table 12). Notably, digits are the only closed-category category to exhibit a positive correlation with class names.
The selective coding data sheds light on this trend. The most frequent digit-related patterns in our dataset fall under Distinguisher × Human-Named Convention and Distinguisher × Locally Specific Concept. These naming strategies are used to distinguish among similar entities (e.g., arg1, arg2, tile3) or to embed system-specific references (e.g., m34, cp437) into variable or type names. Such distinctions are especially useful in parameters and declarations, where there is no syntactic support for disambiguation outside of naming.
In contrast, function-level disambiguation is often handled by the language itself, through overloading, polymorphism, or naming conventions focused on behavior; making digits largely unnecessary or even undesirable in that context. Their absence from function names reflects this shift: digits encode identity, that is, they act as traceability to specific domain concepts, and distinguishing entities with similar names, rather than encoding purpose or behavior.
Taken together, these findings suggest that digits serve primarily as disambiguators or protocol markers rather than communicative devices for expressing behavior. Their presence in class and parameter names signals static or structural variation, while their avoidance in function names underscores a developer preference for meaningful, descriptive action labels over numerical markers.
@@
This strong correlation reflects the behavioral semantics that prepositions convey in identifier names. As detailed in our selective coding, prepositions frequently express directionality, transformation, conditional activation, or event-driven execution; all of which are very function-oriented behaviors, requiring action to be taken. Overall, prepositions help scope, qualify, and clarify a function’s behavior. Their strong correlation with function names supports the interpretation that developers use them deliberately to encode operational semantics directly into the name, especially in contexts involving transformation, control flow, or event handling.
@@
The results of our analysis support the alternative hypothesis: closed-category parts of speech are meaningfully correlated with specific roles and contexts in source code. Prepositions and conjunctions appear more frequently in function names, where they help express behavioral nuances such as guarded actions, type casting, or alternative execution paths. Digits, by contrast, are most commonly found in class names and parameter declarations, where they signal disambiguation, indexing, or versioning; identifiers rooted in identity rather than behavior or purpose.
@@
Figure 3: Per-category Mann-Whitney U test significance across thresholds. Prepositions dominate across thresholds, while conjunctions and digits contribute more variably.
@@
Closed-Category Type Axial Code Domain GitHub URL Language Commit Hash
@@
Prepositions Type Casting / Interpretation Serialization/Deserialization Libraries https://github.com/msgpack/msgpack-c/ C 306d59d52f752388f276aa5244cdd56be3ad75e0
@@
Prepositions Type Casting / Interpretation Serialization/Deserialization Libraries https://github.com/open-source-parsers/jsoncpp.git C++ ca98c98457b1163cca1f7d8db62827c115fec6d1
Prepositions Type Casting / Interpretation Polyglot Interop Tools / Type Bridge Layers https://github.com/swig/swig C++ 4167b0275adc574f86914e42fb5ef9639b2b25a5
@@
Prepositions Type Casting / Interpretation Polyglot Interop Tools / Type Bridge Layers https://github.com/pybind/pybind11 C++ bc4a66dff0464d0c87291b00a3688999fea5bedc
Prepositions Position / Ordering in Time or Space Data Structure / Algorithm Libraries https://github.com/boostorg/container C++ b13c1fe8a4921328c9ecdbdfc90431d73b2060ac
Prepositions Position / Ordering in Time or Space Data Structure / Algorithm Libraries https://github.com/apache/commons-collections Java a3dca96b37f728aa191ca4b077b6604f2e3d2daa
Prepositions Position / Ordering in Time or Space Compiler / Intermediate Representation Tools https://github.com/TinyCC/tinycc C 36ff4f52b518e1043b3aa2f0a466392c08b33978
Prepositions Position / Ordering in Time or Space Compiler / Intermediate Representation Tools https://github.com/rose-compiler/rose C 5c9d8859d38a320d577d7f592a98f4fb762051fc
Determiners Position in Sequence (Temporal / Recent & Upcoming) Job Queues / Schedulers https://github.com/quartz-scheduler/quartz Java df059e2a6c22ad1466476ae246025f31caffa6e5
Determiners Position in Sequence (Temporal / Recent & Upcoming) Job Queues / Schedulers https://github.com/PerMalmberg/libcron C++ aa3d4368d59abb9d51199755df7419194bc3e42e
Determiners Position in Sequence (Temporal / Recent & Upcoming) Parser Generators / Token Stream Libraries https://github.com/antlr/antlr4 Java cb850120378d000e97da38afb852aa3c554b81f5

### v4

@@ 4.2 Manual Process for Annotating Part-of-Speech
Initially, one author (annotator) is assigned annotate each identifier in the CCID with its grammar pattern. The annotator has experience annotating identifiers with part-of-speech (PoS) tags from prior work Newman_GP ; peruma_test_name_gp . The process is as follows: The annotator is given a split (using Spiral HuckaSpiral ) identifier, along with its type, file path, and line number, to facilitate easy identification of the identifier in the original code. The annotator is permitted to examine the source code from which the identifier originated, if necessary. The annotator is asked to additionally identify and correct mistakes made by Spiral. When the annotator is finished, two additional annotators are asked to validate (agree or disagree) with the annotations created by the original annotator. Any disagreements are discussed and fixed, if required. Furthermore, a fourth annotator assigned their own annotations, which are then compared to the original annotator’s work. Again, disagreements are discussed and fixed. An example disagreement is with the identifier where_len, which is a tricky one because ‘where’ is typically an adverb or conjunction. However, in this case ‘where’ is a reference to a void pointer variable called ‘where’ within the code. So ‘where_len’ is the length of the memory this pointer points to, making ‘where’ a noun-adjunct in this case; describing the type of length. Thus, its grammar pattern is NM N. The Fleiss’ Kappa for this process was .916.
We did not expand abbreviations for a couple reasons. The first is that some abbreviations are more meaningful than their expanded terms (e.g., HTTP, IPv4, SSL) due to their frequent use in abbreviated form by the community. The second reason is that abbreviation expansion techniques are not widely available and vary widely in terms of effectiveness on different types of terms newmanabbrev ; Jingxuan:2023 . Therefore, a realistic worst-case scenario for developers and researchers is that no abbreviation-expansion technique is available to use, and their PoS taggers must work in this worst-case scenario. Whenever we recognize one, we do not split domain-term abbreviations (e.g., Spiral will make IPv4 into IPv 4; we corrected this to IPv4). We do this because it is the view of the authors that they should be recognized and appropriately tagged in their abbreviated (i.e., their most common) form.
@@
Determiner Conjunction Preposition numeral
@@ 5 Evaluation of RQ1: What behavioral roles do closed-category terms play in source code identifiers?
Our evaluation aims to establish, through RQ1 and RQ2: 1) how closed-category terms are used to convey differing types of program behavior, 2) the typical grammatical structure of identifiers containing closed-category terms, and 3) how closed-category term distributions differ across programming context, language, and system domains. This first research question investigates the semantic role of closed-category grammatical patterns in identifier naming. We focus on four closed-category part-of-speech types: prepositions, numerals, determiners, and conjunctions. We present our findings by (1) describing each category’s semantic function using axial codes, (2) summarizing behavioral trends via selective coding, and (3) highlighting shared trends through cross-category synthesis.
5.1 Methodology: Manual Process for Behavioral Annotations
We employed an approach inspired by Straussian grounded theory to analyze variable names in source code and their relationship to program behavior. This multi-phase coding process involved four annotators, combining individual annotations with iterative validation and synthesis to construct a theory grounded in observed naming patterns. The sample used in this study is a subset of the CCID described in Section 4. To construct this subset, we took the top 10 most common grammar patterns (Table 7) and collected all identifiers that followed these patterns; randomly selecting the 10th grammar pattern if multiple patterns had the same frequency. These represent the most common names (i.e., from the perspective of grammatical structure) used in the data set. This totals to 618 identifiers.
@@ 5.1 Methodology: Manual Process for Behavioral Annotations
Annotators grouped similar open codes into higher-level axial codes, focusing on patterns where particular grammatical structures consistently aligned with specific behavioral roles. These axial codes captured mid-level abstractions (e.g., State Variables, Event Triggers), and were documented in the spreadsheet alongside notes justifying the grouping where needed. Each annotator’s axial codes were reviewed by a different annotator for validation. This cross-review process involved reading both the open codes and the proposed axial codes, discussing disagreements, and refining the categories until consensus was reached. The Fleiss’ Kappa for this phase was: .971 for numerals, .996 for Determiners, .976 for Prepositions, and 1.0 for Conjunctions.
@@ 5.2 Numerals in Identifiers
Numerals in identifiers act as compact, semantic indicators of structure, ordering, or version. They are also often used to disambiguate entities and encode numeric conventions. Their meaning is typically inferred through domain knowledge, making them powerful, but potentially hard to understand for those without the requisite domain knowledge.
@@
We created a dual-axis framework for interpreting the meaning of numerals, inspired by a single-axis framework we created in prior work on numerals in identifiers peruma_digits . This framework reflects our observation that numerals contribute information in two distinct ways: (1) the role they play within the local context (e.g., indexing, versioning), and (2) the source of meaning they draw from, which is often external to the immediate source code scope (e.g., domain conventions, technical standards). Every numeral in the set has both a role and a source of meaning; they must be combined to fully understand the numeral. We put an ‘x’ between each combination of ‘Role’ and ‘Source of Meaning’ Axial Code.
  • Role: What functional purpose the numeral serves in the identifier.
    • Distinguisher: The numeral differentiates conceptually similar entities, typically to avoid name collision errors from the compiler (e.g., arg1, tile2).
    • Version Identifier: The numeral encodes versioning information such as protocol revisions or data format versions (e.g., http2, v1).
  • Source of Meaning: Where the interpretation of the numeral originates, typically via convention, tooling, or domain-specific logic.
    • Auto-Generated: The numeral is added automatically by tools, compilers, or naming systems to avoid conflicts (e.g., var1_2, jButton3).
    • Human-Named Convention: The numeral’s meaning is primarily derived from ad hoc developer intent and is not more complex than distinguishing entities manually (e.g., str1, feature2).
    • Locally Specific Concept: The numeral conveys project- or context-specific information, often related to coordinate systems, data structures, or memory layouts (e.g., m33 for matrix row 3 col 3).
    • Technology Term / Standard: The numeral is part of a recognized domain-specific label, format, or protocol (e.g., HTTP2, Neo4j).
@@
    Description: This group captures identifiers where numerals encode positional or logical roles based on system-specific conventions, such as grid layout or data structure indexing.
@@
    Description: This group captures identifiers that include numerals as part of standardized or domain-specific naming conventions, often encoding formats or specifications.
@@
    Description: This group captures identifiers where the numeral signals the version number of a protocol, tool, or technology component.
@@
    Description: This group captures identifiers that are automatically suffixed with a numeral to ensure uniqueness, often generated by tools or compilers.
@@
Consider the identifier m34, which appears in the context of a matrix operation. To fully interpret the numeral 3 in this name, we must consider both its Role and its Source of Meaning. Semantically, the numeral serves as a Distinguisher; uniquely identifying this variable apart from its siblings (such as m32 and m31). However, its complete interpretation depends on its Locally Specific Concept source: the developers have an internal convention that 3 refers to the row index, while 4 refers to the column. Without knowing the Source of Meaning, the numbers can only be interpreted as distinguishing one identifier from another; the meaning of the numerals would remain ambiguous. This illustrates how both axes work together—Role tells us what the numeral is doing, while Source of Meaning tells us how to interpret the value.
@@
Numerals serve as semantic compression tools in source code: conveying versioning, layout, ordering, or configuration state using a minimal footprint. Their power lies in the idea that both the identifier’s author and readers share a certain level of domain knowledge, and thus can understand the meaning of the numeral. Whether distinguishing hosts (host1, host2), signaling protocol versions (http2), or denoting matrix dimensions (m33), numerals rely on prior knowledge to be effective, this makes them:
@@
  • Hard to understand when overused without documentation or when the reader lacks background information/experience that the author assumed they would have
Numerals are structural shortcuts in the mental models of developers; a quick way to convey a lot of information in a small number of characters.
@@ 5.3 Prepositions in Identifiers
Prepositions in identifiers express spatial, temporal, or logical relationships. They are the most versatile (i.e., most axial codes) and frequently used closed-class grammatical structure in our dataset. Prepositions typically convey transformation, control conditions, event triggers, source origin, or context membership. Because only a subset of these are dual-axis (Boolean Flow), we inline the definitions with our examples, unlike with numerals, where we separate them.
@@ 5.4 Determiners in Identifiers
    Definition: This group encompasses identifiers that reference a population or subset, typically using quantifiers such as all, any, or some to guide iteration, filtering, or policy logic.
@@ 5.5 Conjunctions in Identifiers
Conjunction-based identifiers are rare but expressive. They signal compound behavior, dual-mode interfaces, or gated logic—often making hidden control flow or semantic relationships visible. Their rarity likely stems from the fact that developers often express conjunctions in logic rather than names. But when used, they highlight either an intent to emphasize control-flow behavior or to capture structural duality within a single name.
@@ 5.6 Cross-Category Synthesis
Across numerals, determiners, prepositions, and conjunctions, developers use closed-class grammatical structures to encode compact, behavior-rich semantics in identifiers. While each PoS category exhibits distinct tendencies, analysis of grammar patterns reveals broader functional themes and stylistic consistencies across categories.
@@
Interestingly, booleans appear in all three of these contexts, but each is a different flavor; a way of expressing behavior that is unique to the closed-category terms used in the boolean identifier.
@@
    Numerals imply procedural differentiation (method1, step2) or timeline indexing when appearing in coordinated identifiers (m31, m32).
@@
    Numerals embed indexing, uniqueness, and domain roles
@@
These layered forms serve as semantic shortcuts to convey complex behavior with minimal words. They compress conditions, transformations, order, and relationships into concise forms that aim to assist program comprehension and understanding without excess verbiage.
@@ 5.7 Summary of RQ1
Through qualitative analysis of closed-category terms in identifiers, we have uncovered and explored how these compact grammatical forms play a central role in expressing program behavior. Each part-of-speech category contributes distinct semantic functions, ranging from transformation and scoping to control flow and logical composition.
@@
Determiner Conjunction Numeral Preposition
@@
We begin by analyzing the distribution of four closed categories: prepositions, determiners, conjunctions, and numerals, across five programming contexts and three programming languages. We discuss which categories are most frequent in which contexts/languages and consider how those patterns may reflect the communicative goals of the developer. We then extend this analysis to system-level domain context, comparing the normalized frequency of closed-category term usage between domain-specific and general-purpose systems.
@@ 6.1 Closed-Category Term Usage Across Programming Contexts, Programming Languages, and System Domains
We now examine how differing contexts and closed-category grammar patterns relate to one another, and whether programming language further conditions their usage. We begin by analyzing cross-language correlations in the usage of closed-category terms, followed by an exploration of correlations in how these terms are used across different program contexts. We provide Table 8, which shows frequencies and percentages for PoS and terms, to help the reader understand what types of terms are most prevalent. However, for this research question, we rely primarily on Tables 9, 10, 11, and 12, which present the results of our Pearson Chi-square tests and standardized Pearson residuals. Using these, we highlight common patterns, terms, and the contexts or languages to which these patterns are correlated.
@@ 6.1.1 Language-Specific Differences in Closed-Category Term Usage
Methodology. To perform our Chi-Square test, we use the CCID described in Section 4.1. We count how many times each closed-category PoS appears in C++, Java, or C code by analyzing all 1,001 identifiers that contain closed-category terms. For example, we might find that there were 20 numerals in our data set found in C++ code, and 5 numerals in Java code. Once we have these frequencies, we apply the Chi-Square test and Standardized Pearson residuals with Bonferroni correction to determine overall significance and per-part-of-speech significance, respectively.
@@ C C++ Java
  • Numerals (D) are modestly underrepresented in Java (residual = –1.65), suggesting a mild tendency to avoid numeric suffixes in Java naming.
@@
While we did not find significant statistical evidence linking closed-category tag usage to programming language, the residual analysis and qualitative trends suggest mild idiomatic differences, particularly around numeral usage and determiner phrasing. For example, in our Axial Code data from RQ1, Population/Subpopulation Reference identifiers were found in Java (21, 50%) and C++ (18, 42%) more than in C(3, 7%). These patterns may reflect broader stylistic conventions or design idioms of each language, but should be interpreted cautiously given the statistical outcome.
@@ 6.1.2 Context-Specific Differences in Closed-Category Term Usage
Methodology. To perform our Chi-Square test, we use the CCID described in Section 4.1. We analyzed all 1,001 identifiers that contained closed-category terms, and counted how many times a closed-category term appears in one of our five code contexts: Attribute, Function, Class, Declaration, or Parameter. Once we have these frequencies, we apply the Chi-Square test and Standardized Pearson residuals with Bonferroni correction to determine overall significance and per-part-of-speech significance, respectively.
@@
The selective coding data from RQ1 explains this pattern. Conjunction-based grammar patterns tend to express compound logic, dual-purpose behavior, or guarded activation, which are most relevant when naming behaviors or actions rather than static values. For example, Guarded Action / Conditional Enablement patterns such as load_if_needed or activate_if_enabled appear in function names to encode preconditions or gating logic directly into the identifier.
@@
While we did find a significant correlation between conjunctions and functions, it is essential to note that our dataset contains 50 conjunctions, meaning that while we have identified potential trends, further research on a larger sample is likely necessary.
@@
Numerals (D). Closed-category grammar patterns that include numerals often feature numerals such as 1, 2, 0, 3, and 4, as shown in Table 8. Numerals are significantly positively correlated with parameter names and class names, and significantly negatively correlated with function names (Standardized Pearson residual = 4.72 for class names, 4.25 for parameters, and –5.51 for functions; Table 12). Notably, numerals are the only closed-category category to exhibit a positive correlation with class names.
The selective coding data sheds light on this trend. The most frequent numeral-related patterns in our dataset fall under Distinguisher × Human-Named Convention and Distinguisher × Locally Specific Concept. These naming strategies are used to distinguish among similar entities (e.g., arg1, arg2, tile3) or to embed system-specific references (e.g., m34, cp437) into variable or type names. Such distinctions are especially useful in parameters and declarations, where there is no syntactic support for disambiguation outside of naming.
In contrast, function-level disambiguation is often handled by the language itself, through overloading, polymorphism, or naming conventions focused on behavior; making numerals largely unnecessary or even undesirable in that context. Their absence from function names reflects this shift: numerals encode identity, that is, they serve as a means of traceability to specific domain concepts and distinguish entities with similar names, rather than encoding purpose or behavior.
Taken together, these findings suggest that numerals serve primarily as disambiguators or protocol markers rather than communicative devices for expressing behavior. Their presence in class and parameter names signals static or structural variation, while their avoidance in function names underscores a developer’s preference for meaningful, descriptive action labels over numerical markers.
@@
This strong correlation reflects the behavioral semantics that prepositions convey in identifier names. As detailed in our selective coding, prepositions frequently express directionality, transformation, conditional activation, or event-driven execution; all of which are highly function-oriented behaviors, requiring action to be taken.
@@
The results of our analysis support the alternative hypothesis: closed-category parts of speech are meaningfully correlated with specific roles and contexts in source code. Prepositions and conjunctions appear more frequently in function names, where they help express behavioral nuances such as guarded actions, type casting, or alternative execution paths. Numerals, by contrast, are most commonly found in class names and parameter declarations, where they signal disambiguation, indexing, or versioning; identifiers rooted in identity rather than behavior.
@@
Figure 3: Per-category Mann-Whitney U test significance across thresholds. Prepositions dominate across thresholds, while conjunctions and numerals contribute more variably.
@@
Closed-Category Type Axial Code Domain GitHub URL Language
@@
Prepositions Type Casting / Interpretation Serialization/Deserialization Libraries https://github.com/msgpack/msgpack-c/ C
@@
Prepositions Type Casting / Interpretation Serialization/Deserialization Libraries https://github.com/open-source-parsers/jsoncpp.git C++
@@
Prepositions Type Casting / Interpretation Polyglot Interop Tools / Type Bridge Layers https://github.com/pybind/pybind11 C++
Prepositions Type Casting / Interpretation Polyglot Interop Tools / Type Bridge Layers https://github.com/swig/swig C++